A school organized a book fair and in this book fair a book seller is selling his books under the following rules:


There are three different packages available.


First package contains 2 Islamic books, 2 Science books and 2 Geography books, second package contains 2 Islamic books, 4 Science books and 1 Geography books and third package contains 3 Islamic books, 4 Science books and 5 Geography books. The book fair has a total of 250 Islamic books, 300 Science books, and 270 Geography books. First package makes a profit of Rs. 120, second package makes Rs.100 and third package makes Rs.270 per pack.


How many packs should be made to maximize book fair profits?


What will the profit be?

1
Expert's answer
2021-12-27T16:11:40-0500

X1 = First package

X2 = Second package

X3 = Third package

Objective Function:

"Max Z =120X_1+100X_2+270X_3"

Subject to:

"2X_1+2X_2+3X_3=250"

"2X_1+4X_2+4X_3=300"

"2X_1+1X_2+5X_3=270"

"X_1,X_2,X_3>0"


The result of LPP is as follows:



As per above LPP output,

"X_1=82"

"X_2=16"

"X_3=18"


Then, profit will be:

"Max Z =(120\times82)+(100\times16)+(270\times18)=16300"


This means 82 first packages, 16 second packages and 18 third packages packs should be made to maximize book fair profit.

In this case, the expected book fair profit would be Rs. 16,300.